OpenStudy (harrisonshark):

The hypotenuse of a right triangle is 25 cm, and the shorter leg is 15 cm. Find the length of the other leg.

OpenStudy (anonymous):

sqrt10...

OpenStudy (harrisonshark):

butiful

OpenStudy (anonymous):

simply use pthagoras thoremm...

OpenStudy (harrisonshark):

ahhhhh

jigglypuff314 (jigglypuff314):

use the pythagorean theorem c^2 = a^2 + b^2 where c is the hypotenuse and a and b are the legs if c = 25 and a = 15 then 25^2 = 15^2 + b^2 solve for b